OpenStudy (anonymous):

(b) A wheel of radius 5 in. is rotating 45°/sec. What is the linear speed v, the angular speed in RPM and the angular speed in rad/sec?

OpenStudy (anonymous):

\[v=r \omega \] This is the formula to go from translational velocity to angular velocity

OpenStudy (anonymous):

\[\frac{ 45rev }{ 1\sec }*\frac{ 60\sec }{ 1\min }=2700 rev/min\]

OpenStudy (anonymous):

That was the angular speed

OpenStudy (anonymous):

1 revolution = 2pi \[2700\frac{ rev }{ \min }*\frac{ 2\pi rads }{ 1rev }=5400\pi \frac{ rads }{ \min }\]

OpenStudy (anonymous):

Its 45 degrees per second do you just assume that it is a rev

OpenStudy (anonymous):

Oh sorry. I didn't notice that

OpenStudy (anonymous):

change degrees to radians

OpenStudy (anonymous):

45 degrees= pi/4 radians btw

OpenStudy (anonymous):

Of so then what does that do for the rev/min

OpenStudy (anonymous):

\[\frac{ 45^{o} }{ 1\sec }*\frac{ 1rev }{ 360^{o} }=\frac{ 1rev }{ 8\sec }\]

OpenStudy (anonymous):

\[\frac{ 1rev }{ 8\sec }*\frac{ 60\sec }{ 1\min }=7.5\frac{ rev }{ \min }\]